OpenStudy (anonymous):

Please help Consider the leading term of the polynomial function. What is the end behavior of the graph? Describe the end behavior and provide the leading term. -3x5 + 9x4 + 5x3 + 3

OpenStudy (anonymous):

because a) the degree is odd, and b) the leading coefficient is negative, this mean \[\lim_{x\to -\infty}f(x)=\infty\] and \[\lim_{x\to \infty}f(x)=-\infty\]